solution A 30% acid, solution B is 100% acid. How many liters should be combined to make 28 liters of a solution that is 50%acid?
X + Y = 28
30% 100%=50%
30x 100y 50% of 24 liters <

Answer by Alan3354(30993)

solution A 30% acid, solution B is 100% acid. How many liters should be combined to make 28 liters of a solution that is 50%acid?
X + Y = 28
30% 100%=50%
30x 100y 50% of 24 liters < ------------------
x = amount of solution A
y = amount of B
---------------
x + y = 28 --> y = 28-x
0.3x + y = 0.5*28 = 14
--------------
y = 28-x
0.3x + 28-x = 14
-0.7x = -14
x = 20 liters
y = 8 liters
-----------------
I would have used A and B instead of x and y, but the answer is the same.